# Proceedings of the ASME 2025International Mechanical Engineering Congress & Exposition (IMECE)

### v1

@@ 2.1 Problem description
The computational flow domain of the ejector is illustrated in Fig. 2. The cylindrical coordinate system defines $x$ as the axial direction, $r$ as the radial direction, and $\theta$ as the azimuthal direction. This longitudinal slice through the ejector centerline shows the motive inlet (red, $x/d=-13$), the suction inlet (green, $-11\leq x/d\leq-8$), diffuser outlet (black, $x/d=23$), and bounding walls (cyan). The corresponding velocity vector components are defined by $u_{x}$, $u_{r}$, $u_{\theta}$. The origin is placed at the center of the motive throat. A converging needle controls the inlet opening. The inlet flow is annular. Table 1 reports the 3D simulations’ boundary conditions.
@@
Motive inlet $P_{m}=63.5$ bar $\dot{m}_{m}=57.4$ g/s $T_{m}=296$ K $Y_{\text{CO}_{2}(g)}=0$ $\mathcal{H}=0.38$ mm $P_{m}=63.5$ bar $\nabla{\vec{u}}=\vec{0}$ $T_{m}=296$ K $Y_{\text{CO}_{2}(g)}=0$ $\mathcal{H}=0.31$ mm
@@ 2.3 Numerics
We performed 3D transient simulations with a variable time step algorithm allowing minimum and maximum time step sizes of 10 ns and 1 ms, respectively, limiting the change in time step size to 25% of the previous time step size as the solution advances [19]. The pressure-velocity coupling is achieved using the Pressure Implicit Splitting Order (PISO) algorithm [21], the pressure Poisson equation is discretized with a HYPRE Bi-Conjugate Stabilized (BiCGSTAB) scheme [22], and the discretized momentum and species transport equations are solved with the Successive Over Relaxation (SOR) scheme with an over-relaxation factor of 1.0. The time step size is calculated based on convective, diffusive, and compressibility stability limits according to a Courant-Friedrichs-Lewy (CFL) number. Specifically, for convection $\text{CFL}_{c}=\vec{u}\Delta t/\Delta$ and we require $\text{CFL}_{c}\leq 0.5$. Meanwhile,he CFL-like number for diffusion is $\text{CFL}_{d}=(\nu+\nu_{\text{T}})\Delta t/\Delta^{2}\leq 0.25$, and the CFL-like number for compressibility is $\text{CFL}_{a}=a\Delta t/\Delta\leq 5$. From the three constraints, the most restrictive constraint is enforced in the simulations [19]. Here, $\nu$ is the local momentum diffusivity, $\nu_{\text{T}}$ is the local turbulent momentum diffusivity, $\Delta$ is the local mesh cell edge length, $a$ is the local speed of sound, and $\Delta t$ is the global time step size. The variable time step algorithm is based on the grid size controlled by Adaptive Mesh Refinement (AMR), the number of iterations at the previous time step, and the CFL numbers.
@@
Таблица 2: Details of the grid implemented with $\Delta_{\text{base}}=4.0$ mm and adaptive mesh refinement
Region $n$ $\Delta$ ($\mu$m) Refinement condition
@@ 3 Validation and Verification
The numerical model from the previous section was validated against measured data from the experimental facility at the Ray W. Herrick Laboratories [5]. Table 1 summarizes the boundary conditions in the motive region, where Dirichlet-type pressure, temperature, and Neumann-type velocity boundary conditions are provided and the simulation solves for the mass flow rate. Similarly, for the suction region, a Dirichlet-type mass flow rate, temperature, and Neumann-type pressure boundary conditions are provided, and the code solves for the pressure.
@@ 4 Internal Flow Topology
A sharp radial density gradient can be observed at $x/d=2$ in Fig. 5(b) across the jet interface, decreasing $\bar{\rho}_{m}/\bar{\rho}_{s}$ from 2.75 to 1.25 within a radial distance of $\Delta r/d=0.5$. The radial distance, $\Delta r/d$, of the radial density gradient can be observed to increase along the axial direction; however, the differential ratio $\Delta\bar{\rho}_{m}/\bar{\rho}_{s}\rightarrow 0$. This observation clearly shows a radially expanding but axisymmetric mixing layer along the shear layer and a decaying density of the motive jet along the axial direction. Furthermore, until $x/d\leq 4$ at $-0.25\leq r/d\leq 0.25$ a slight decrease in the mixture density can be observed, and a slight increase when $r/d\rightarrow 1$. These minor changes result from local heat transfer along the wall, which does not majorly affect the flow field in R3.
@@
A prominent potential core can be observed in R1, R2, and R3, where our previous instantaneous resolved scale model shows that the shear layer is not near-wall [14]. However, the jet in R4 transformed into a wall-bounded turbulent flow, decaying the potential core completely by the wall shear effect as shown in Fig. 6(a). The rate of decay of potential core due to the wall-shear effect is well described by the fit
@@ 5 Similarity Analysis of the Flow Field in Regime 3
Previous discussion in section 4 shows that the flow is axisymmetric ($\partial/\partial\theta=0$), isothermal ($\nabla\mu=0$), and assuming a quasi-steady condition ($\partial/\partial t=0$) for simplicity, the conservation and transport equations (5) and (5) can be modified as follows:
@@
$\phi=\tilde{\phi}+\phi^{{}^{\prime\prime}},\qquad\text{where}\quad\tilde{\phi}=\frac{\overline{\rho_{m}\phi}}{\bar{\rho}_{m}}$ (10)
We have also given the relation to the Reynolds averaged quantities (denoted by overbars). Observe that $\overline{\rho_{m}\phi^{{}^{\prime\prime}}}=0$ but $\ \overline{\phi^{{}^{\prime\prime}}}\neq 0$.
Applying the Favre-averaging (10) to the mass and momentum conservation equations (8) and (5), we obtain
@@
$\qquad-\frac{1}{\bar{\rho}_{m}}\frac{\partial}{\partial X}(\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{x}^{{}^{\prime\prime}}})-\frac{1}{r\bar{\rho}_{m}}\frac{\partial}{\partial r}(r\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{r}^{{}^{\prime\prime}}})$ (12)
@@
$\nu\left[\frac{1}{3}\frac{\partial^{2}\overline{u_{x}^{{}^{\prime\prime}}}}{\partial X^{2}}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial\overline{u_{x}^{{}^{\prime\prime}}}}{\partial r}\right)\right]$
was in deriving Eq. (5) because it is negligible for large Reynolds number flow [26]. In Eq.5, the terms $(1/\bar{\rho}_{m})(\partial/\partial X)(\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{x}^{{}^{\prime\prime}}})$ and $(1/r\bar{\rho}_{m})(\partial/\partial r)(r\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{r}^{{}^{\prime\prime}}})$ represents the Reynolds stress terms in axial and radial directions, which are the typical unclosed terms modeled in different turbulence models [15].
Attempting a similarity transformation of the flow field, $\tilde{u}_{x},\tilde{u}_{r},[\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{x}^{{}^{\prime\prime}}}+(\tilde{p}-\tilde{p}_{0})]$, and $(\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{r}^{{}^{\prime\prime}}})$ are represented as a function of the Favre-averaged centerline velocity, $\tilde{u}_{c}$, and shape functions $f(\eta)$, $g(\eta)$, $h(\eta)$, and $k(\eta)$, respectively. Here, $p_{0}$ is a constant pressure, and $\eta$ is a transformed coordinate combining the axial $X$ and radial $r$ coordinates. Based on the discussion in section 4 (recall Fig. 6(b)), the centerline velocity is represented as
@@
The mathematical representations of the transformed variables are is then as follows,
@@
$\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{x}^{{}^{\prime\prime}}}+(\tilde{p}-\tilde{p}_{0})=\tilde{u}_{c}^{2}\bar{\rho}_{m}h(\eta)$ (14d)
$\overline{\rho_{m}u_{x}^{{}^{\prime\prime}}u_{r}^{{}^{\prime\prime}}}=\tilde{u}_{c}^{2}\bar{\rho}_{m}k(\eta)$ (14e)
Applying the transformations (14) to the mass and momentum conservation equations (11) and (5) and using Eqs. (13), we obtain
$g=\frac{2}{\left(\frac{1}{\eta}+\frac{X}{\bar{\rho}_{m}}\frac{\partial\bar{\rho}_{m}}{\partial r}\right)}$
×$\phantom{g=}\times\int_{0}^{\infty}\left(\frac{1}{\eta}+\frac{X}{\bar{\rho}_{m}}\frac{\partial\bar{\rho}_{m}}{\partial r}\right)\left(\eta\frac{df}{d\eta}-\frac{X}{\bar{\rho}_{m}}f\frac{\partial\bar{\rho}_{m}}{\partial X}-nf\right)d\eta$ (15)
@@
$-\frac{X}{\bar{\rho}_{m}}\left(h\frac{\partial\bar{\rho}_{m}}{\partial X}+\frac{\partial\bar{\rho}_{m}}{\partial r}\right)$
$+\frac{1}{\text{Re}_{c}}\left[-\frac{n(n-1)f}{3}+\frac{2}{3}\left(1-n+\frac{3}{2\eta^{2}}\right)\eta\frac{df}{d\eta}+\frac{(\eta^{2}-3)}{3}\frac{d^{2}f}{d\eta^{2}}\right]$ (16)
@@
It is clear from Eqs. (5) and (5) that the similarity transformation of the conservation equations presents incomplete similarity because $\bar{\rho}_{m}$, which is defined in terms of $\tanh(X,r)$ and cannot be expressed as a sole function of $\eta$, remains in the equations. However, at $X/d\geq 7$, $\nabla\bar{\rho}_{m}\rightarrow\vec{0}$ can be transformed into a completely self-similar flow field.
Nevertheless, we can represent the analytical fit for density, i.e., Eq. (2), in terms of a new dimensionless similarity variable (namely, $\varepsilon=X/d$) along with transformed coordinate $\eta=r/X$, specifically:
$\bar{\rho}_{m}(\varepsilon,\eta\varepsilon)=a\left[\tanh\left(2\eta\varepsilon b+c\right)+\tanh\left(-2\eta\varepsilon b+c\right)\right]+e$ (17)
@@
$\frac{X}{\bar{\rho}_{m}}\left(h\frac{\partial\bar{\rho}_{m}}{\partial X}+\frac{\partial\bar{\rho}_{m}}{\partial r}\right)\\ =-\frac{4.918\varepsilon h}{\bar{\rho}_{m}}\left[\tanh\left(2\eta\varepsilon b+c\right)+\tanh\left(-2\eta\varepsilon b+c\right)\right]\\ -\frac{\varepsilon h}{\bar{\rho}_{m}}\left(1.6374\varepsilon+13.3152\right)\\ +\frac{\varepsilon dha}{\bar{\rho}_{m}}\Bigg{[}2\eta\left(-0.5277\varepsilon+b\right)\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)\\ -2\eta b\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)-2\eta(-0.5277\varepsilon+b)\operatorname{sech}^{2}\left(-\eta\varepsilon b+c\right)\\ +2\eta b\operatorname{sech}^{2}\left(-\eta\varepsilon b+c\right)\\ +\frac{2\left(0.0681\varepsilon-1.0086\right)}{d}\left[\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)+\operatorname{sech}^{2}\left(-2\eta\varepsilon b+c\right)\right]\Bigg{]}\\ +\frac{2\varepsilon ab}{\bar{\rho}_{m}}\left[\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)-\operatorname{sech}^{2}\left(-\eta\varepsilon b+c\right)\right]$ (18)
@@
[…]
@@
$G=2/\left[\eta^{-1}+X\bar{\rho}_{m}^{-1}(\partial\bar{\rho}_{m}/\partial r)\right]=2/\left(\eta^{-1}+F\left[\operatorname{tanh(\varepsilon,\eta)}\right]\right)$
and $\lim_{\eta\rightarrow\infty}G=2$, which is a constant. The first term in Eq. (19) is related to the boundary conditions ($f|_{\eta\rightarrow\infty}-f|_{\eta=0}$) and the third integral is related to the velocity distribution. However, all other integrals in Eq. (19) are a function of density gradient, which leads to incompleteness in the self-similarity reduction requiring further care. Its nature is similar to $\tanh(\varepsilon,\eta)$ using asymptotic approximation. Furthermore, the presence of $\tanh(\varepsilon,\eta)$ in the expression for $g(\eta)$ also influences the transformed momentum equation Eq. (5). Indeed, the effect of the density gradient is confirmed by observing the radial velocity distribution in Fig. 7.

### v2

@@ 2.1 Problem description
The computational flow domain of the ejector is illustrated in Fig. 2. The cylindrical coordinate system defines $x$ as the axial direction, $r$ as the radial direction, and $\theta$ as the azimuthal direction. This longitudinal slice through the ejector centerline shows the motive inlet (red, x/d=13x/d=-13), the suction inlet (green, $11\leq x/d\leq-8$), diffuser outlet (black, $x/d=23$), and bounding walls (cyan). The corresponding velocity vector components are defined by $u_{x}$, $u_{r}$, $u_{\theta}$. The origin is placed at the center of the motive throat. A converging needle controls the inlet opening. The inlet flow is annular. Table 1 reports the 3D simulations’ boundary conditions.
@@
Motive inlet $P_{m}=63.5$ bar $\dot{m}_{m}=57.4$ g/s $T_{m}=296$ K $Y_{\text{CO}_{2}(g)}=0$ $\mathcal{H}=0.38$ $m$ $P_{m}=63.5$ bar $\nabla{\vec{u}}=\vec{0}$ $T_{m}=296$ K $Y_{\text{CO}_{2}(g)}=0$ $\mathcal{H}=0.31$ $m$
@@ 2.3 Numerics
We performed 3D transient simulations with a variable time step algorithm allowing minimum and maximum time step sizes of 10 ns and 1 ms, respectively, limiting the change in time step size to 25% of the previous time step size as the solution advances [19]. The pressure-velocity coupling is achieved using the Pressure Implicit Splitting Order (PISO) algorithm [21], the pressure Poisson equation is discretized with a HYPRE Bi-Conjugate Stabilized (BiCGSTAB) scheme [22], and the discretized momentum and species transport equations are solved with the Successive Over Relaxation (SOR) scheme with an over-relaxation factor of 1.0. The time step size is calculated based on convective, diffusive, and compressibility stability limits according to a Courant-Friedrichs-Lewy (CFL) number. Specifically, for convection $\text{CFL}_{c}=|\vec{u}|\Delta t/\Delta$ and we require $\text{CFL}_{c}\leq 0.5$. Meanwhile, the CFL-like number for diffusion is $\text{CFL}_{d}=(\nu+\nu_{\text{T}})\Delta t/\Delta^{2}\leq 0.25$, and the CFL-like number for compressibility is $\text{CFL}_{a}=a\Delta t/\Delta\leq 5$. From the three constraints, the most restrictive constraint is enforced in the simulations [19]. Here, $\nu$ is the local momentum diffusivity, $\nu_{\text{T}}$ is the local turbulent momentum diffusivity, $\Delta$ is the local mesh cell edge length, $a$ is the local speed of sound, and $\Delta t$ is the global time step size. The variable time step algorithm is based on the grid size controlled by Adaptive Mesh Refinement (AMR), the number of iterations at the previous time step, and the CFL numbers.
The flow domain is discretized with a Cartesian grid generated using a cut cell method implemented in the CONVERGE CFD software [19]. The grid is a Cartesian grid with uniform edge lengths in the three coordinate directions. The baseline grid, $\Delta_{\text{base}}$, for this study has $\Delta_{\text{base}}=4$ mm grid resolution based on our grid convergence study [14]. To avoid cell defeaturing and loss of orthogonality at constricted geometry, the base grid is refined at these locations to a refinement level of $n=5$, minimizing local grid size as $\Delta=\Delta_{\text{base}}/2^{n}$ to 0.125mm0.125\ $\text${mm} [5]. AMR is also implemented, which depends on the gradients as listed in Table 2, adapting the computational mesh to the evolving features of the flow and species fields.
Таблица 2: Details of the grid implemented with $\Delta_{\text{base}}=4.0$ $m$ and adaptive mesh refinement
@@ 3 Validation and Verification
The numerical model from the previous section was validated against measured data from the experimental facility at the Ray W. Herrick Laboratories [5]. Table 1 summarizes the boundary conditions in the motive region, where Dirichlet-type pressure, temperature, and Neumann-type velocity boundary conditions are provided, and the simulation solves for the mass flow rate. Similarly, for the suction region, a Dirichlet-type mass flow rate, temperature, and Neumann-type pressure boundary conditions are provided, and the code solves for the pressure.
@@ 4 Internal Flow Topology
A sharp radial density gradient can be observed at $x/d=2$ in Fig. 5(b) across the jet interface, decreasing $\bar{\rho}_{m}/\bar{\rho}_{s}$ from 2.75 to 1.25 within a radial distance of $\Delta r/d=0.5$. The radial distance, $\Delta r/d$, of the radial density gradient can be observed to increase along the axial direction; however, the differential ratio $\Delta\bar{\rho}_{m}/\bar{\rho}_{s}\rightarrow 0$. This observation clearly shows a radially expanding but axisymmetric mixing layer along the shear layer and a decaying density of the motive jet along the axial direction. Furthermore, until $x/d\leq 4$ at $0.25\leq r/d\leq 0.25$ a slight decrease in the mixture density can be observed, and a slight increase when $r/d\rightarrow 1$. These minor changes result from local heat transfer along the wall, which does not significantly affect the flow field in R3.
@@
A prominent potential core can be observed in R1, R2, and R3, where our previous instantaneous resolved scale model shows that the shear layer is not near-wall [14]. However, the jet in R4 transformed into a wall-bounded turbulent flow, decaying the potential core completely by the wall shear effect as shown in Fig. 6(a). The rate of decay of the potential core due to the wall-shear effect is well described by the fit
@@ 5 Similarity Analysis of the Flow Field in Regime 3
Previous discussion in section 4 shows that the flow is axisymmetric ($\partial/\partial\theta=0$), isothermal ($\nabla\mu=0$), and assuming a quasi-steady condition ($\partial/\partial t=0$) for simplicity, the conservation and transport equations (Appendix) and (Appendix) can be modified as follows:
@@
$\phi=\tilde{\phi}+\phi^{\prime\prime},\qquad\text{where}\quad\tilde{\phi}=\frac{\overline{\rho_{m}\phi}}{\bar{\rho}_{m}}$ (10)
[…]
Applying the Favre-averaging (24) to the mass and momentum conservation equations (22) and (Appendix), we obtain
@@
$\qquad-\frac{1}{\bar{\rho}_{m}}\frac{\partial}{\partial X}(\overline{\rho_{m}u_{x}^{\prime\prime}u_{x}^{\prime\prime}})-\frac{1}{r\bar{\rho}_{m}}\frac{\partial}{\partial r}(r\overline{\rho_{m}u_{x}^{\prime\prime}u_{r}^{\prime\prime}})$ (12)
@@
$\nu\left[\frac{1}{3}\frac{\partial^{2}\overline{u_{x}^{\prime\prime}}}{\partial X^{2}}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial\overline{u_{x}^{\prime\prime}}}{\partial r}\right)\right]$
was in deriving Eq. (Appendix) because it is negligible for large Reynolds number flow [26]. In Eq.Appendix, the terms $(1/\bar{\rho}_{m})(\partial/\partial X)(\overline{\rho_{m}u_{x}^{\prime\prime}u_{x}^{\prime\prime}})$ and $(1/r\bar{\rho}_{m})(\partial/\partial r)(r\overline{\rho_{m}u_{x}^{\prime\prime}u_{r}^{\prime\prime}})$ represents the Reynolds stress terms in axial and radial directions, which are the typical unclosed terms modeled in different turbulence models [15].
Attempting a similarity transformation of the flow field, $\tilde{u}_{x},\tilde{u}_{r},[\overline{\rho_{m}u_{x}^{\prime\prime}u_{x}^{\prime\prime}}+(\tilde{p}-\tilde{p}_{0})]$, and $(\overline{\rho_{m}u_{x}^{\prime\prime}u_{r}^{\prime\prime}})$ are represented as a function of the Favre-averaged centerline velocity, $\tilde{u}_{c}$, and shape functions $f(\eta)$, $g(\eta)$, $h(\eta)$, and $k(\eta)$, respectively. Here, $p_{0}$ is a constant pressure, and $\eta$ is a transformed coordinate combining the axial $X$ and radial $r$ coordinates. Based on the discussion in section 4 (recall Fig. 6(b)), the centerline velocity is represented as
@@
The mathematical representations of the transformed variables are then as follows,
@@
$\overline{\rho_{m}u_{x}^{\prime\prime}u_{x}^{\prime\prime}}+(\tilde{p}-\tilde{p}_{0})=\tilde{u}_{c}^{2}\bar{\rho}_{m}h(\eta)$ (14d)
$\overline{\rho_{m}u_{x}^{\prime\prime}u_{r}^{\prime\prime}}=\tilde{u}_{c}^{2}\bar{\rho}_{m}k(\eta)$ (14e)
Applying the transformations (29) to the mass and momentum conservation equations (27) and (Appendix) and using Eqs. (30), we obtain
$g=\frac{1}{\exp\left[\int_{0}^{\eta}\left(\frac{1}{\eta}+\frac{X}{\bar{\rho}_{m}}\frac{\partial\bar{\rho}_{m}}{\partial r}\right)d\eta\right]}$
$\phantom{g=}\times\int_{0}^{\infty}\exp\left[\int_{0}^{\eta}\left(\frac{1}{\eta}+\frac{X}{\bar{\rho}_{m}}\frac{\partial\bar{\rho}_{m}}{\partial r}\right)d\eta\right]$
[…]
@@
$-\frac{X}{\bar{\rho}_{m}}\left(h\frac{\partial\bar{\rho}_{m}}{\partial X}+k\frac{\partial\bar{\rho}_{m}}{\partial r}\right)$
$+\frac{1}{\text{Re}_{c}}\left[-\frac{n(1-n)f}{3}+\frac{2}{3}\left(1-n+\frac{3}{2\eta^{2}}\right)\eta\frac{df}{d\eta}+\frac{(\eta^{2}+3)}{3}\frac{d^{2}f}{d\eta^{2}}\right]$ (16)
@@
It is clear from Eqs. (5) and (5) that the similarity transformation of the conservation equations presents incomplete similarity because $\bar{\rho}_{m}$, which is defined in terms of $\tanh(X,r)$ cannot be expressed as a sole function of $\eta$, remains in the equations. However, at $X/d\geq 7$, $\nabla\bar{\rho}_{m}\rightarrow\vec{0}$ can be transformed into a completely self-similar flow field only when (i) n=1n=-1 (to eliminate the dependence of $\text{Re}_{c}$ on $X$) or (ii) $n=1$ while simultaneously assuming that $\text{Re}_{c}\rightarrow\infty$.
Either way, we can represent the analytical fit for density, i.e., Eq. (2), in terms of a new dimensionless similarity variable (namely, $\varepsilon=X/d$) along with transformed coordinate $\eta=r/X$, specifically:
$\bar{\rho}_{m}(\varepsilon,\eta)=a\left[\tanh\left(2\eta\varepsilon b+c\right)+\tanh\left(-2\eta\varepsilon b+c\right)\right]+e$ (17)
@@
$\frac{X}{\bar{\rho}_{m}}\left(h\frac{\partial\bar{\rho}_{m}}{\partial X}+k\frac{\partial\bar{\rho}_{m}}{\partial r}\right)=\frac{h}{a}\left(1-\frac{e}{\overline{\rho}_{m}}\right)\varepsilon\frac{da}{d\varepsilon}\\ +\frac{2a}{\overline{\rho}_{m}}\varepsilon\left(\eta h\varepsilon\frac{db}{d\varepsilon}+bk\right)\left[\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)-\operatorname{sech}^{2}\left(-2\eta\varepsilon b+c\right)\right]\\ +\frac{ah}{\overline{\rho}_{m}}\varepsilon\frac{dc}{d\varepsilon}\left[\operatorname{sech}^{2}\left(2\eta\varepsilon b+c\right)+\operatorname{sech}^{2}\left(-2\eta\varepsilon b+c\right)\right]+\frac{h}{\overline{\rho}_{m}}\varepsilon\frac{de}{d\varepsilon}$ (18)
@@
$g=\frac{1}{G}\int_{0}^{\infty}G\left(\eta\frac{df}{d\eta}-\frac{\varepsilon d}{\overline{\rho}_{m}}fF\left[\operatorname{tanh}(\varepsilon,\eta)\right]-nf\right)d\eta$ (19)
@@
$G=\left[\eta+\int\exp\left(\frac{\varepsilon d}{\overline{\rho}_{m}}F\left[\operatorname{tanh}(\varepsilon,\eta)\right]\right)d\eta\right]$
and $\lim_{\eta\rightarrow\infty}(G/\eta)=1+\exp(\varepsilon d/\overline{\rho}_{m})$, which is a constant. The second integral is a function of the density gradient, which leads to incompleteness in the self-similarity reduction and requires further attention. Its nature is similar to $\tanh(\varepsilon,\eta)$ using asymptotic approximation. Furthermore, the presence of $\tanh(\varepsilon,\eta)$ in the expression for $g(\eta)$ also influences the transformed momentum equation Eq. (5). Indeed, the effect of the density gradient is confirmed by observing the radial velocity distribution in Fig. 7.